# A Dual Adaptive Assignment Approach for Robust Graph-Based Clustering

### v1

@@ Abstract.
Graph clustering is an essential aspect of network analysis that involves grouping nodes into separate clusters. Recent developments in deep learning have resulted in advanced deep graph clustering techniques, which have proven effective in many applications. Nonetheless, these methods often encounter difficulties when dealing with the complexities of real-world graphs, particularly in the presence of noisy edges. Additionally, many denoising graph clustering strategies tend to suffer from lower performance compared to their non-denoised counterparts, training instability, and challenges in scaling to large datasets. To tackle these issues, we introduce a new framework called the Dual Adaptive Assignment Approach for Robust Graph-Based Clustering (RDSA). RDSA consists of three key components: (i) a node embedding module that effectively integrates the graph’s topological features and node attributes; (ii) a structure-based soft assignment module that improves graph modularity by utilizing an affinity matrix for node assignments; and (iii) a node-based soft assignment module that identifies community landmarks and refines node assignments to enhance the model’s robustness. We assess RDSA on various real-world datasets, demonstrating its superior performance relative to existing state-of-the-art methods. Our findings indicate that RDSA provides robust clustering across different graph types, excelling in clustering effectiveness and robustness, including adaptability to noise, stability, and scalability.
@@ 1. INTRODUCTION
Graph clustering is crucial in network analysis, spanning physics, bioinformatics, and social sciences. Real-world graphs naturally exhibit clusters, representing sets of nodes with similar characteristics. These clusters hold significant implications, delineating social groups in friendship networks (Bedi and Sharma, 2016), functional modules within protein-interaction networks (Chen and Yuan, 2006), and thematic groups of academic papers within citation networks (Sen et al., 2008). Recently, breakthroughs in deep learning have led to a paradigm shift in artificial intelligence and machine learning, achieving great success in many important tasks, including graph clustering. The main techniques used in deep graph clustering are MLP-based graph clustering (Gao et al., 2019; Jia et al., 2019; Cui et al., 2020; Liu et al., 2023; Pal and Mitra, 1992), GNN-based graph clustering (Wang et al., 2017; Park et al., 2019; Fan et al., 2020; Cheng et al., 2021; Xia et al., 2021; Wang et al., 2021), and hybrid-based graph clustering (Bo et al., 2020; Peng et al., 2021; Tu et al., 2021; Hassani and Khasahmadi, 2020; Zhao et al., 2021; Liu et al., 2022b, a; Yang et al., 2023a, b, 2022).
@@ 4.2. Structure-based Soft Assignment
Due to the randomized initial state of the framework, the training process may become unstable and converge to a local minimum, resulting in suboptimal outcomes and reduced robustness. To enhance robustness in structure-based soft assignment, we employ an auxiliary information loss to assist this assignments, inspired by DGCluster (Bhowmick et al., 2024). Let $\boldsymbol{M}\in\boldsymbol{V}$ represent a subset of nodes, and $\boldsymbol{N}$ denote the pairwise membership matrix of $\boldsymbol{M}$, where $\boldsymbol{N}(i,j)=1$ if nodes $i$ and $j$ belong to the same cluster, and $\boldsymbol{N}(i,j)=0$ otherwise. The auxiliary information loss is defined as:
@@
However, some datasets lack auxiliary information. In these cases, clusters generally exhibit higher structural density in their central regions. For datasets without auxiliary information, these central regions can serve as substitutes for the missing auxiliary data. In many cases, the reliability of auxiliary information is uncertain, so we assign it a minor weight, denoted as $\alpha$, to maintain equilibrium with the modularity maximization loss. The total loss function of the structure-based soft assignment can be formulated as:
@@ 4.3. Node-based Soft Assignment
In the proposed framework, we previously introduced a structure-based soft assignment method that focuses on modularity maximization. However, this approach only considers the graph structure. To make better use of node information, we now include a node-based soft assignment. SLIM (Zhu et al., 2022) is a graph pooling method that identifies several landmarks in the graph to represent it as a whole. Unlike clustering centroids, these landmarks are updated dynamically during training. In our framework, we select landmarks to represent each community and assign nodes to these landmarks using a soft assignment approach. This node-based soft assignment aims to maximize relationships within clusters while minimizing relationships between clusters. Thus, we emphasize stabilizing these relationships to improve the framework’s robustness by enhancing the connectivity between nodes and landmarks within each community.
The first step in the node-based soft assignment is selecting landmarks, which involves choosing the $k$ most representative nodes in each community. Since landmarks are crucial for community representation, they need to be highly connected within their communities. Modularity serves as a measure of this connectivity. Based on the modularity scores from the previous step, we select the $k$ nodes with the highest scores within each community to ensure that each landmark accurately represents its community. This selection process allows landmarks to better capture the local structure and nuances compared to centroids and makes them more robust to outliers, as their positions are less affected by extreme values. The landmark selection process can be formulated as follows:
@@ 6.1. Datasets
We evaluate our proposed method on two medium-scale datasets (Cora (Sen et al., 2008) and Citeseer (Sen et al., 2008)), three large-scale datasets (Pubmed (Sen et al., 2008), Amazon Photo (Shchur et al., 2019), and Amazon Computers (Shchur et al., 2019)), and one extra large scale dataset (ogbn arxiv (Hu et al., 2020)). The statistics of these datasets are summarized in Table 1.
@@ 6.3. Baselines Comparison
In this section, we undertake experiments to illustrate the effectiveness of RDSA using eight baseline models across five datasets. Specifically, the clustering algorithms under comparison can be categorized into three groups: MLP-based techniques (AGE (Cui et al., 2020), SCGC (Liu et al., 2023)), GNN-based approaches (DCRN (Liu et al., 2022a), ARGA (Pan et al., 2020)), and hybrid methodologies (SDCN (Bo et al., 2020), SUBLIME (Liu et al., 2022b), AGC-DRR (Gong et al., 2022), CONVERT (Yang et al., 2023b)).
The experimental findings are summarised in Table 2. Herein, we delineate several vital observations: 1) Our method surpasses existing high-performance graph clustering techniques, achieving state-of-the-art results; 2) Hybrid approaches exhibit superior performance compared to MLP-based and GNN-based methods across most datasets, suggesting a potential inadequacy in considering diverse aspects of graph data in the latter methods; 3) Our method, benefiting from dual soft assignment objectives that comprehensively leverage both structural and attribute information of graph data, outperforms other hybrid methods. Our approach showcases superior performance across $85\%$ of evaluation metrics across diverse datasets. As illustrated with the citation network datasets, our method achieves a $4.96\%$ improvement in ACC, a $6.00\%$ improvement in NMI, a $9.00\%$ improvement in ARI, and a $1.89\%$ improvement in F1 score compared to the second-best performing method. Meanwhile, our method demonstrates superior performance in large-scale datasets such as Amazon photos and computers, confirming our method’s scalability. These results underscore the efficacy of our proposed technique in enhancing clustering performance within deep graph clustering frameworks. Additionally, scatter plots of the embedding space post dimensionality reduction are provided in Figure 2, illustrating our method’s ability to discern distinct clusters within the embedded space effectively.
@@ 7. CONCLUSION
In this paper, we propose a novel robust deep graph clustering method, RDSA, which integrates structure based soft assignment, node based soft assignment, and attribute embedding to enhance the model’s adaptability to noisy data, stability, and scalability. Extensive experiments on six benchmark datasets, along with ablation studies, demonstrate the superior performance of our method compared to state-of-the-art approaches. These findings underscore its efficacy in handling noisy data, maintaining stability, and scaling to large datasets, while also validating the significance of each model component in achieving this performance. Future work will focus on extending our method to a broader range of graph-related tasks, such as node classification and link prediction, to further validate its versatility and robustness. Additionally, we plan to explore its application in real-world problems like recommendation systems, where graph-based representations play a crucial role, and molecular research, where accurate clustering of molecular structures is essential. These extensions will allow us to assess the practical impact of RDSA in diverse and complex domains.

### v4

# RDSA: A Robust Deep Graph Clustering Framework via Dual Soft Assignment

@@ Abstract.
Graph clustering is an essential aspect of network analysis that involves grouping nodes into separate clusters. Recent developments in deep learning have resulted in graph clustering, which has proven effective in many applications. Nonetheless, these methods often encounter difficulties when dealing with real-world graphs, particularly in the presence of noisy edges. Additionally, many denoising graph clustering methods tend to suffer from lower performance, training instability, and challenges in scaling to large datasets compared to non-denoised models. To tackle these issues, we introduce a new framework called the Robust Deep Graph Clustering Framework via Dual Soft Assignment (RDSA). RDSA consists of three key components: (i) a node embedding module that effectively integrates the graph’s topological features and node attributes; (ii) a structure-based soft assignment module that improves graph modularity by utilizing an affinity matrix for node assignments; and (iii) a node-based soft assignment module that identifies community landmarks and refines node assignments to enhance the model’s robustness. We assess RDSA on various real-world datasets, demonstrating its superior performance relative to existing state-of-the-art methods. Our findings indicate that RDSA provides robust clustering across different graph types, excelling in clustering effectiveness and robustness, including adaptability to noise, stability, and scalability. The code is available at https://github.com/EsiksonX/RDSA.
@@ 1. INTRODUCTION
Graph clustering is crucial in network analysis, spanning physics, bioinformatics (Wu et al., 2024), and social sciences (Kong et al., 2024). Real-world graphs naturally exhibit clusters, representing sets of nodes with similar characteristics. These clusters hold significant implications, delineating social groups in friendship networks (Bedi and Sharma, 2016), functional modules within protein-interaction networks (Chen and Yuan, 2006), and thematic groups of academic papers within citation networks (Sen et al., 2008). Recently, breakthroughs in deep learning have led to a paradigm shift in artificial intelligence and machine learning, achieving great success in many important tasks, including graph clustering. The main techniques used in deep graph clustering are MLP-based graph clustering (Gao et al., 2019; Jia et al., 2019; Cui et al., 2020; Liu et al., 2023; Pal and Mitra, 1992), GNN-based graph clustering (Wang et al., 2017; Park et al., 2019; Fan et al., 2020; Cheng et al., 2021; Xia et al., 2021; Wang et al., 2021), and hybrid-based graph clustering (Bo et al., 2020; Peng et al., 2021; Tu et al., 2021; Hassani and Khasahmadi, 2020; Zhao et al., 2021; Liu et al., 2022b, a; Yang et al., 2023a, b, 2022).
@@ 4.2. Structure-based Soft Assignment
The random initial state in clustering tasks can lead to an unstable training process, causing convergence to a local minimum. To mitigate this problem, we use an auxiliary information loss, inspired by DGCluster (Bhowmick et al., 2024), to guide the model toward global optimal convergence and improve stability. Let $\boldsymbol{M}\in\boldsymbol{V}$ represent a subset of nodes, and $\boldsymbol{N}$ denote the pairwise membership matrix of $\boldsymbol{M}$, where $\boldsymbol{N}(i,j)=1$ if nodes $i$ and $j$ belong to the same cluster, and $\boldsymbol{N}(i,j)=0$ otherwise. The auxiliary information loss is defined as:
@@
However, some datasets lack additional information. In these cases, clusters tend to show higher structural density in their central regions. For datasets without this information, these central regions can act as substitutes for the missing data. Let $\alpha$ represent the weight of auxiliary information loss. Often, the reliability of auxiliary information is uncertain. When we are unsure about the confidence of this information, we typically assign a lower weight to mitigate the effects of this uncertainty (Bhowmick et al., 2024). The total loss function for the structure-based soft assignment can be formulated as:
@@ 4.3. Node-based Soft Assignment
In the proposed framework, we introduce a structure-based soft assignment method focused on modularity maximization. However, this primarily considers only the graph structure. Inspired by SLIM (Zhu et al., 2022), which identifies representative nodes (landmarks) for each community to characterize the entire graph, The landmarks are dynamically updated during training, which is different from traditional fixed centroids. Community landmarks effectively capture the distribution of communities and improve the ability to distinguish between different communities, which means that landmarks can provide strong discriminative power in unsupervised clustering. Therefore, landmarks can be introduced to improve the robustness of graph clustering.
The node-based soft assignment process begins with the selection of landmarks. Specifically, the $k$ most representative nodes within each community are chosen based on connectivity, assessed via modularity. As landmarks are central to representing community structures, they are chosen to be well-connected within their respective communities to provide a stable representation. By selecting nodes with the highest modularity scores, each landmark effectively captures the local community structure and nuances, performing more robustly than centroids and being less affected by outliers. This landmark selection process can be formalized as follows:
@@ 6.1. Datasets
We evaluate our proposed method on two medium-scale datasets (Cora and Citeseer), three large-scale datasets (Pubmed, Amazon Photo, and Amazon Computers), and two extra large scale dataset (ogbn arxiv, and ogbn products ). The statistics of these datasets are summarized in Table 1. .
@@ 6.3. Baselines Comparison
In this section, we present experiments to evaluate the performance of RDSA compared with eight baseline models across seven datasets. Table LABEL:table:performance provides a summary of the experimental results. The primary findings are as follows: 1) Our method outperforms existing graph clustering techniques, setting new performance benchmarks; 2) Hybrid approaches generally exceed the performance of both MLP-based and GNN-based methods on most datasets, suggesting that MLP and GNN models may not effectively capture the diversity within graph data; 3) RDSA uses dual soft assignment objectives to incorporate both structural and attribute information from graph data, leading to improved performance over other hybrid methods. Specifically, our approach achieves the top scores in $85\%$ of evaluation metrics across the tested datasets. For instance, on the citation network datasets, our method provides a $4.96\%$ increase in ACC, a $6.00\%$ improvement in NMI, a $9.00\%$ increase in ARI, and a $1.89\%$ increase in F1 score over the second-best method, supporting the effectiveness of our framework.
Additionally, we performed experiments on two large-scale datasets. While all baseline methods, except MAGI, encounter memory limitations on these datasets, our method consistently delivers the best performance. On the ogbn-arxiv dataset, RDSA surpasses MAGI by $8.99\%$ in accuracy and $8.12\%$ in ARI, while on the ogbn-products dataset, RDSA exceeds MAGI by $11.5\%$ in accuracy and $32.9\%$ in ARI, indicating RDSA’s higher consistency in clustering boundaries. Although minor differences in information alignment and small-cluster identification result in NMI and F1 scores that are slightly lower than those of MAGI, these variations do not affect the overall clustering quality. These results underscore the efficacy of our proposed framework in enhancing clustering performance within deep graph clustering frameworks.
@@ 7. CONCLUSION
In this paper, we propose a novel robust deep graph clustering method, RDSA, which integrates structure based soft assignment, node based soft assignment, and attribute embedding to enhance the model’s adaptability to noisy data, stability, and scalability. Extensive experiments on seven benchmark datasets, along with ablation studies, demonstrate the superior performance of our method compared to state-of-the-art approaches. These findings underscore its efficacy in handling noisy data, maintaining stability, and scaling to large datasets, while also validating the significance of each model component in achieving this performance. Future work will focus on extending our method to a broader range of graph-related tasks, such as node classification and link prediction, to further validate its versatility and robustness. Additionally, we plan to explore its application in real-world problems like recommendation systems, where graph-based representations play a crucial role, and molecular research, where accurate clustering of molecular structures is essential. These extensions will allow us to assess the practical impact of RDSA in diverse and complex domains.
Acknowledgements.
This work is partially funded by the XJTLU Research Development Fund [Grant Number RDF-21-02-050] and the Young Data Scientist Program of the China National Astronomical Data Center [Grant Number NADC2023YDS08].